# SAQ: Pushing the Limits of Vector Quantization through Code Adjustment and Dimension Segmentation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
@@ 1. Introduction
With the proliferation of machine learning, embedding models (Devlin et al., 2019; Radford et al., 2021; Shvetsova et al., 2022; Li et al., 2022) are widely used to map diverse data objects, including images, videos, texts, e-commerce goods, genes, and proteins, to high-dimensional vector embeddings that encode their semantic information. A core operation on these vectors is nearest neighbor search (NNS(Elastic, [n.d.]; pgvector, [n.d.]; Wang et al., 2021; SingleStore, [n.d.]; Qdrant, [n.d.]; Yang et al., 2020; Mohoney et al., 2023), which retrieves vectors that are the most similar to a query vector from a vector dataset. NNS underpins many important applications including content search (e.g., for images and videos), recommender systems, bio-informatics, and retrieval-argumented generation (RAG) for LLMs (Lewis et al., 2020). However, exact NNS requires a linear scan in high-dimensional space, rendering it impractical for large-scale datasets. To address this, approximate NNS (ANNS) has been widely adopted (Huang et al., 2020; Xiong et al., 2020), which trades exactness for efficiency by returning most of the top-$k$ nearest neighbors. Many vector indexes (Malkov and Yashunin, 2018a; Yin et al., 2025) and vector databases (Elastic, [n.d.]; pgvector, [n.d.]; Wang et al., 2021; SingleStore, [n.d.]; Qdrant, [n.d.]) support ANNS as the core functionality.
@@ 2.3. Extended RaBitQ
The normalization and projection of E-RaBitQ are the same as RaBitQ but E-RaBitQ use uses the following codebook $\mathcal{G}_{r}$ to quantize $\mathbf{o}_{p}$
@@ 3.1. Quantization Procedure
1 Input: round limit $r$, data vector $\mathbf{o}$, start point $\bar{\mathbf{o}}$, value $v_{\max}$
@@ 4.1. Motivation and Problem Formulation
In the reverse direction, if we concentrate the data vectors’ variances into certain dimensions, we can remove the dimensions with negligible variances. PCA is widely used for this purpose. A recent study (Yang et al., 2025) indicates that the high-dimensional vector coordinates, once rotated by a PCA matrix, exhibit a long-tailed variance distribution, which implies that only a few dimensions hold most of the variance (Figure 5(b)). This characteristic was used to create an approximate distance estimator that uses the leading dimensions. These techniques are commonly known as dimension reduction. Dimension reduction is less general, as it is highly dependent on the data distribution. When PCA fails to polarize the variance, a fixed dimension reduction ratio will result in poor accuracy.
@@ 4.2. Quantization Plan Construction
According to our experiments and the analysis of RaBitQ (Gao et al., 2024), the relative error of quantization a segment (or vector) is proportional to $2^{-B}$, where $B$ is the number of bits used to quantify the segment. This is natural because with 1 more bit, we can reduces the quantization resolution by $1/2$. Moreover, we observe that, after PCA projection, the value distribution of a dimension almost follows a normal distribution with a mean close to 0, as illustrated in Figure 7(b). Therefore, we propose the following expression to model the error introduced in the index phase:
@@
1 Input: vector dimension $D$, total bit quota $Q_{quota}$
@@ 4.3. Multi-Stage Distance Estimation
According to our observation, after PCA projection, the value distribution of a dimension almost follows a normal distribution with a mean close to 0, as illustrated in Figure 7(b). The variance of expression above can be written as:
@@ 5. EXPERIMENTAL Evaluation
    PQ (Jégou et al., 2011): A popular quantization method that is generally used with a high compression rate and is widely used in industry. We use the Faiss PQ implementation (Faiss, [n.d.]). PQ supports two settings $nbits=4$ and $nbits=8$, which is the number of bits used to represent each sub-vector after quantization. According to  (Gao et al., 2024), $nbits=8$ produces consistently better than $nbits=4$, so we report the result under this setting.
@@
We do not compare with OPQ (Ge et al., 2013), LSQ (Martinez et al., 2016), and scalar quantization (SQ) because they are outperformed by RaBitQ. To test the performance of the vector quantization methods for ANNS, we build IVF index (Jegou et al., 2010) for all datasets following RaBitQ. In particular, IVF groups the vectors into clusters and scans the top-ranking clusters for each query. Following the common setting of IVF, we used 4,096 clusters for the datasets. We are aware that proximity graph indexes (Malkov and Yashunin, 2018b; Fu et al., 2017) are also popular, but combining with vector indexes is not our focus and we leave it to future work. All methods were optimized with the SIMD instructions with AVX512 to ensure a fair efficiency comparison, and the performance improvements of SAQ generalize across platforms.
@@
Datasets. We used four public real-world datasets that have various dimensionalities and data types, as shown in Table 2. These datasets include widely adopted benchmarks for the evaluation of ANN algorithms(Aumüller et al., 2020; Li et al., 2020; Gao and Long, 2024; Gao et al., 2024; Yang et al., 2025) (DEEP and GIST [^4] https://www.cse.cuhk.edu.hk/systems/hash/gqr/datasets.html) and embeddings generated by language models. The MSMARCO [^5] https://huggingface.co/datasets/Cohere/msmarco-v2.1-embed-english-v3 dataset contains the embeddings for the TREC-RAG Corpus 2024 (TREC-RAG, [n.d.]) embedded with the Cohere Embed V3 English model (Cohere, [n.d.]) and the OpenAI-1536666https://huggingface.co/datasets/Qdrant/dbpedia-entities-openai3-text-embedding-3-large-1536-1M dataset is produced by model text-embedding-3-large of OpenAI (OpenAI, [n.d.]). We used the query sets provided by the DEEP and GIST datasets. For MSMARCO and OpenAI-1536, we randomly removed 1,000 vectors from each dataset and used them as query vectors.

### v2

@@ Abstract.
conference: ACM on Management of Data; May 31–June 5, 2026; Bengaluru, Indiacopyright: ccjournalyear: 2025journalvolume: 3journalnumber: 6 (SIGMOD)article: 359publicationmonth: 12doi: 10.1145/3769824
Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/howarlii/saq/.
@@ 1. Introduction
With the proliferation of machine learning, embedding models (Devlin et al., 2019; Radford et al., 2021; Shvetsova et al., 2022; Li et al., 2022) are widely used to map diverse data objects, including images, videos, texts, e-commerce goods, genes, and proteins, to high-dimensional vector embeddings that encode their semantic information. A core operation on these vectors is nearest neighbor search (NNS(Elastic, ; pgvector, ; Wang et al., 2021; SingleStore, ; Qdrant, ; Yang et al., 2020; Mohoney et al., 2023), which retrieves vectors that are the most similar to a query vector from a vector dataset. NNS underpins many important applications including content search (e.g., for images and videos), recommender systems, bio-informatics, and retrieval-argumented generation (RAG) for LLMs (Lewis et al., 2020). However, exact NNS requires a linear scan in high-dimensional space, rendering it impractical for large-scale datasets. To address this, approximate NNS (ANNS) has been widely adopted (Huang et al., 2020; Xiong et al., 2020), which trades exactness for efficiency by returning most of the top-$k$ nearest neighbors. Many vector indexes (Malkov and Yashunin, 2018b; Yin et al., 2025) and vector databases (Elastic, ; pgvector, ; Wang et al., 2021; SingleStore, ; Qdrant, ) support ANNS as the core functionality.
@@ 2.3. Extended RaBitQ
The normalization and projection of E-RaBitQ are the same as RaBitQ but E-RaBitQ uses the following codebook $\mathcal{G}_{r}$ to quantize $\mathbf{o}_{p}$
@@ 3.1. Quantization Procedure
1Input: round limit $r$, data vector $\mathbf{o}$, start point $\bar{\mathbf{o}}$, value $v_{\max}$
@@ 4.1. Motivation and Problem Formulation
In the reverse direction, if we concentrate the data vectors’ variances into certain dimensions, we can remove the dimensions with negligible variances. PCA is widely used for this purpose. A recent study (Yang et al., 2025) indicates that the high-dimensional vector coordinates, once rotated by a PCA matrix, exhibit a long-tailed variance distribution, which implies that only a few dimensions hold most of the variance (Figure 5). This characteristic was used to create an approximate distance estimator that uses the leading dimensions. These techniques are commonly known as dimension reduction. Dimension reduction is less general, as it is highly dependent on the data distribution. When PCA fails to polarize the variance, a fixed dimension reduction ratio will result in poor accuracy.
@@ 4.2. Quantization Plan Construction
According to our experiments and the analysis of RaBitQ (Gao et al., 2024), the relative error of quantization a segment (or vector) is proportional to $2^{-B}$, where $B$ is the number of bits used to quantify the segment. This is natural because with 1 more bit, we can reduces the quantization resolution by $1/2$. Moreover, we observe that, after PCA projection, the value distribution of a dimension almost follows a normal distribution with a mean close to 0, as illustrated in Figure 7. Therefore, we propose the following expression to model the error introduced in the index phase:
@@
1Input: vector dimension $D$, total bit quota $Q_{quota}$
@@ 4.3. Multi-Stage Distance Estimation
According to our observation, after PCA projection, the value distribution of a dimension almost follows a normal distribution with a mean close to 0, as illustrated in Figure 7. The variance of expression above can be written as:
@@ 5. EXPERIMENTAL Evaluation
    PQ (Jégou et al., 2011): A popular quantization method that is generally used with a high compression rate and is widely used in industry. We use the Faiss PQ implementation (Faiss, ). PQ supports two settings $nbits=4$ and $nbits=8$, which is the number of bits used to represent each sub-vector after quantization. According to  (Gao et al., 2024), $nbits=8$ produces consistently better than $nbits=4$, so we report the result under this setting.
@@
We do not compare with OPQ (Ge et al., 2013), LSQ (Martinez et al., 2016), and scalar quantization (SQ) because they are outperformed by RaBitQ. To test the performance of the vector quantization methods for ANNS, we build IVF index (Jegou et al., 2010) for all datasets following RaBitQ. In particular, IVF groups the vectors into clusters and scans the top-ranking clusters for each query. Following the common setting of IVF, we used 4,096 clusters for the datasets. We are aware that proximity graph indexes (Malkov and Yashunin, 2018a; Fu et al., 2017) are also popular, but combining with vector indexes is not our focus and we leave it to future work. All methods were optimized with the SIMD instructions with AVX512 to ensure a fair efficiency comparison, and the performance improvements of SAQ generalize across platforms.
@@
Datasets. We used four public real-world datasets that have various dimensionalities and data types, as shown in Table 2. These datasets include widely adopted benchmarks for the evaluation of ANN algorithms(Aumüller et al., 2020; Li et al., 2020; Gao and Long, 2024; Gao et al., 2024; Yang et al., 2025) (DEEP and GIST [^4] https://www.cse.cuhk.edu.hk/systems/hash/gqr/datasets.html) and embeddings generated by language models. The MSMARCO [^5] https://huggingface.co/datasets/Cohere/msmarco-v2.1-embed-english-v3 dataset contains the embeddings for the TREC-RAG Corpus 2024 (TREC-RAG, ) embedded with the Cohere Embed V3 English model (Cohere, ) and the OpenAI-1536666https://huggingface.co/datasets/Qdrant/dbpedia-entities-openai3-text-embedding-3-large-1536-1M dataset is produced by model text-embedding-3-large of OpenAI (OpenAI, ). We used the query sets provided by the DEEP and GIST datasets. For MSMARCO and OpenAI-1536, we randomly removed 1,000 vectors from each dataset and used them as query vectors.